Graph in the rectangular coordinate systems provided.
3x = 15

Answer by Edwin McCravy(20060)

Graph in the rectangular coordinate systems provided.
3x = 15


Notice that the equation

3x = 15

has no term in the variable y, but only in x.  And yet 
we are to graph it on a coordinate system that has both
an x-axis and a y-axis, we need to add the term " 0y " to
either side of the equation so it will have a place to
substitute values for y.  So I'll pick the left side to
add " 0y " to.  So we write it this way:

3x + 0y = 15

Now let's get some points.

Choose an arbitrary value for y and solve for x.

Suppose we let y be, say, 7, then

  3x + 0y = 15 becomes
3x + 0(7) = 15
   3x + 0 = 15
       3x = 15
        x = 5

So one point is (x,y) = (5,7)

Choose another arbitrary value for y and solve for x.

Suppose this time we let y be, say, -4, then

   3x + 0y = 15 becomes
3x + 0(-4) = 15
    3x + 0 = 15
        3x = 15
         x = 5

So another point is (x,y) = (5,-4)

Plot those two points 



Now draw a straight line through them:



Notice that the line is VERTICAL.  This will always be the case
when there is no y-term in the original equation.  You can just
memorize this fact and save time graphing it because you won't 
need to bother finding two points.

(Incidentally it is also true that if variable x is missing in the
equation, the graph will be a HORIZONTAL line.)

Edwin